OpenStudy (anonymous):

x+1divided by x-1 + 1over x

OpenStudy (lgbasallote):

you must mean \(\Large \frac{x+1}{x-1} + \frac{1}{x}\) right?

OpenStudy (callisto):

\[\frac{x+1}{x-1} + \frac{1}{x}\]\[=\frac{(x+1)x}{(x-1)x} + \frac{1(x-1)}{x(x-1)}\]\[=\frac{(x^2+x)}{(x-1)x} + \frac{(x-1)}{x(x-1)}\]\[=\frac{(x^2+x)+(x-1)}{x(x-1)}\]\[=\frac{(x^2+2x-1)}{x(x-1)}\]